(6 points) Find a vector orthogonal to both (-5, -4,0) and to (0, -4, –5) of the form (1,1 -
Answer #1

<1, a, b>.<-5,-4,0>=0 » - $-4a=0 => a = -ş 4. <la, b>.<o,-4,-$>=0 - 4 a-şb=0 - 4(-5) = $b b=1 Required vector, <1 D>